OpenStudy (anonymous):

f(x) = −4x3 + 8x2 + 21x Find the zeros of the polynomial

OpenStudy (rsadhvika):

start by factoring out the gcf

OpenStudy (anonymous):

2

OpenStudy (rsadhvika):

2 is not a factor of 21 check again and wat about x ?

OpenStudy (jack1):

\[f(x) = −4x^3 + 8x^2 + 21x\] first factor the equation so x is a factor of everything in the above equation, so factor that out first \[f(x) = −4x^3 + 8x^2 + 21x \rightarrow \rightarrow \rightarrow x(−4x^2 + 8x + 21x)\]

OpenStudy (jack1):

sorry, becomes : \[f(x) = x(−4x^2+8x+21)\]

OpenStudy (jack1):

\[f(x) = x(−4x^2+8x+21)\] further factors down to...? factors of 21 = 1 and 21 or 7 and 3 factors of -4x^2 are 2x and -2x or 4x and -x

OpenStudy (jack1):

which combination would add to 8x...? so 7 and 3 and 2x and -2x in \[f(x) = x(−4x^2+8x+21)\] \[f(x) = -x(2x - 7)(2x + 3)\]

OpenStudy (jack1):

so your zeros are at f(x) = 0 \[f(x) = -x(2x - 7)(2x+3)\] \[ 0 = -x(2x - 7)(2x+3)\] so call the brackets A and B now (bear with me here) there are 3 options for the zeros either x = 0 [as 0 times ( A ) times ( B ) ]= 0 or A = zero {[as x times ( 0 ) times ( B ) ]= 0} or B = zero {[as x times ( A ) times ( 0 ) ]= 0}

OpenStudy (jack1):

so the zeros are at: x = 0 [as 0 times ( A ) times ( B ) ]= 0 or A = zero {[as x times ( 0 ) times ( B ) ]= 0} so if A = 0, (2x -7) =0 2x - 7 = 0 2x = 7 x = 7/2 = 3.5 or B = zero {[as x times ( A ) times ( 0 ) ]= 0} so if A = 0, (2x +3) =0 2x +3 = 0 2x = -3 x = -3/2 = -1.5

OpenStudy (jack1):

...sorry it's so long winded, but does that all make sense dude?